# Towards Data-Driven Metrics for Social Robot Navigation Benchmarking

### v1

@@ V Experimental Results
We propose a Gated Recurrent Unit-based neural network (GRU [3]) ALT metric model to demonstrate the feasibility of data-driven approaches and to serve as a baseline in future work. This section describes the data transformations applied, the architecture used, the training procedure, as well as quantitative and qualitative results.
@@ V-A Data transformations
  • Features indicating potential danger to humans: minimum time to collision, maximum cost of fear, and maximum cost of panic as defined in [17].
@@
The contextual values in $c_{t}$ are computed from the free-form text describing the context of the task using queries to an LLM (Anthropic Claude). The prompt used was: “I will give a task description for a robot. I want you to reply with the percentile (a number from 0 to 100) that corresponds to <VARIABLE> in comparison with that of other tasks you could imagine. I don’t want an explanation, only the percentile. Take your time to think, but respond with a single integer from 0 to 100.”, where <VARIABLE> is substituted with each of the strings in Table II. The output of the LLM is later normalised to the range 0-1.
@@ V-B Training procedure
After discarding the trajectories used as control questions, we create a 0.9/0.05/0.05 train/validation/test split and train a GRU-based neural network. The network consists of a GRU connected to an MLP. The GRU has 4 layers with 256-unit hidden layers, and the output of the last step of its sequence is fed into an MLP with a 256-unit hidden layer and a 1-unit output layer. The hidden layers use LeakyReLU activation with a $0.01$ negative slope, and the output layer uses sigmoid activation. The model is trained with a batch size of 32 and a learning rate of 1e-5, using MSE loss with the target of the raters’ scores. The training follows an early-stopping strategy with a patience of 20 epochs. Checkpoint saving is performed every time the loss of the validation set improves.
V-C Results
Training resulted in an MSE loss of $0.0469$ for the validation dataset, achieved in epoch $208$. The final test MSE loss is $\mathbf{0.0477}$, equivalent to a MAE of $\mathbf{0.162}$. The MSE and MAE of the model with respect to the mean of the control questions were $0.0062$ and $0.066$, respectively. To avoid over-representing these data points, control questions were not included in the dataset. Fig. 5 shows the mean scores for the control questions, along with their standard deviations, and the estimation the model made for each of them.
@@ V-C Results
To test the model from a qualitative perspective, we generated three simulated scenarios in which people are located frontally between the robot’s initial position and the goal position (Fig. 6(a)). The robot moves at different speeds following several trajectories (a total of $101$) that deviate a variable distance from the straight line to the goal. The top left image in Fig. 6(a) shows a subset of these trajectories as dotted lines. Additionally, four different contexts are considered for each scenario, each with distinct urgency, importance, and risk among other variables.
(a) Robot trajectories and scenarios of Figs. 6(b), 6(c), and 6(d).
@@
Figure 6: This figure shows the output of the learned metric for different trajectory variations, situations, contexts and speeds. In all scenarios, the robot and the goal are located in the same position, as indicated in Fig 6(a), in orange at the bottom and green at the top, respectively. In Figs. 6(b) and 6(d), there is a single human; in the first case static, in the second case approaching the location where the robot starts. In Fig.6(c), there are three humans, as depicted in Fig.6(a). To get to its goal the robot follows different trajectories, with varying degrees of divergence from the central line (top-left image of Fig. 6(a)). The contexts are as shown on the top-left of each figure: (lab) is for “A robot is working with lab samples. The samples contain a deadly virus”; (fire) is for “A restaurant robot is looking for a fire extinguisher, as it just detected a fire”; (office) is for “An office assistant robot keeps track of who is in the office today”, and (fragile) is for “A delivery robot is navigating in a hospital. It works with fragile objects”. Speed is color-coded:   0.20 m/s   0.40 m/s   0.80 m/s   1.60 m/s.
Figures 6(b), 6(c), and 6(d) show the metric model output for these three scenarios. The horizontal axis represents the maximum deviation of the trajectory from the straight line. The sign refers to the side, left or right, where the deviation occurs. The plots show the score assigned by the trained model to each trajectory.
As these figures show, the model learns to distinguish among contexts. The highest speed is penalized more in the contexts labeled as lab and fragile in Fig. 6, where the robot would generally be expected to move slowly due to the risk of the task. In these two contexts, the robot is also expected to maintain a wider distance from humans, which explains why trajectories closer to the human (i.e., with smaller deviations) receive lower scores than in the other contexts. In contrast, the fire context reflects a high-urgency situation: the lowest speed results in scores close to zero on all trajectories and scenarios, and closer proximity to humans is more acceptable.
@@
Finally, notable differences can be observed between the two scenarios with one person (Figs. 6(b) and 6(d)). When the person moves, the robot should move faster than when the person is static. This aligns with the expectation that the robot should actively maneuver in time to avoid the moving person safely. Additionally, deviating farther from the straight line is more penalized than in the static scenario, which can be explained by the fact that, once the robot has safely avoided the human, the distance from them is greater than in the stationary person scenario, as the person moves away from the robot.
@@ VI Conclusions and Future Work
This paper described how All-encompassing Learned Task-wise metrics [4] can be learned following a data-driven approach in the context of social robot navigation. The experimental results support the idea that the approach is feasible; the quantitative results achieve MSE and MAE of $0.0477$ and $0.162$, respectively. The qualitative results show that the learned metric is sensitive to both the content of the scenarios and the context of the tasks.
@@
Our envisioned future work spans two different areas: a) increasing the dataset size, and b) development of domain-specific deep learning architectures that can more efficiently leverage geometric relationships.

### v2

@@ V Experimental Results
We propose a Gated Recurrent Unit-based neural network (GRU [cho2014learning]) ALT metric model to demonstrate the feasibility of data-driven approaches and to serve as a baseline in future work. This section describes the data transformations applied, the architecture used, the training procedure, and the quantitative and qualitative results.
@@ V-A Data transformations
  • Features indicating potential danger to humans: minimum time to collision, maximum cost of fear, and maximum cost of panic as defined in [singamaneni2023towards].
@@
The contextual values in $c_{t}$ are computed from the free-form text describing the context of the task using queries to an LLM (Anthropic’s Claude 3.7 Sonnet v20250219) [^3] The model was trained using the quantization generated by Claude 3.7 Sonnet v20250219, Prometheus v2.0, gpt-oss-20b, LLaMa 3.3 and Deepseek R1. Claude produced the best model loss.. The prompt used was: “I will give a task description for a robot. I want you to reply with the percentile (a number from 0 to 100) that corresponds to <VARIABLE> in comparison with that of other tasks you could imagine. I don’t want an explanation, only the percentile. Take your time to think, but respond with a single integer from 0 to 100.”, where <VARIABLE> is substituted with each of the strings in Table II. The output of the LLM is later normalized to the range 0-1.
@@ V-B Training procedure
After discarding the trajectories used as control questions, we create a $0.9/0.05/0.05$ train/validation/test split and train a GRU-based neural network. This split was chosen to maximize the amount of data available for training, given the diversity of contexts and trajectories. Considering the dataset’s moderate size, it ensures stable learning while keeping the validation and test sets sufficient to assess generalization. In total, these subsets correspond to $181$ and $211$ trajectories, $44$ and $50$ minutes worth of trajectory data, respectively, which we consider sufficient for evaluation.
The network consists of a GRU connected to an MLP. The GRU has 4 layers with 256-unit hidden layers, and the output of the last step of its sequence is fed into an MLP with a 256-unit hidden layer and a 1-unit output layer. The hidden layers use LeakyReLU activation with a $0.01$ negative slope, and the output layer uses sigmoid activation. The model is trained with a batch size of $32$ and a learning rate of 1e51e-5, using MSE loss with the target of the raters’ scores. The training follows an early-stopping strategy with a patience of $20$ epochs.
V-C Quantitative results
The final test MSE loss is $\mathbf{0.0457}$, equivalent to a MAE of $\mathbf{0.160}$. The MSE and MAE of the model w.r.t. the mean of the control questions were $0.0053$ and $0.053$, respectively. To avoid over-representing these data points, control questions were not included in the dataset. Fig. 3 shows the mean scores for the control questions, along with their standard deviations, and the estimation the model made for them.
@@ V-C Quantitative results
We used Pearson (r) and Spearman (p) correlations to measure how common analytic metrics and the learned metric proposed align with human scores. The results, in Table III, show a much higher alignment between human scores and the proposed ALT metric. Metrics related to proxemics (social space intrusions ratio, and maximum number of near humans) show a certain level of alignment, but it can be considered low-moderate.
r p
success 0.366 0.280
min. distance to humans 0.009 0.397
min. distance to objects -0.006 0.100
collisions with humans -0.385 -0.374
collisions with objects -0.387 -0.365
social space intrusions ratio (A) -0.316 -0.422
social space intrusions ratio (B) -0.271 -0.341
social space intrusions ratio (C) -0.184 -0.252
max. num. of near humans (A) -0.415 -0.420
max. num. of near humans (B) -0.303 -0.328
max. num. of near humans (C) -0.253 -0.283
min. time to collision 0.121 0.313
max. fear -0.341 -0.303
maximum panic -0.128 -0.156
path efficiency ratio 0.105 0.050
cumulative heading changes 0.049 0.035
learned metric (ours) 0.797 0.684
TABLE III: Pearson (r) and Spearman (p) correlations with humans’ scores for different SocNav metrics and the learned one on the test set. (A), (B) and (C) refer to different distance thresholds: $0.4$, $0.6$, and $0.8$. Results yielding p-values greater than $0.05$ have been considered to have low statistical significance and marked in red.
To analyze the influence of different groups of input features on the performance of the model, an ablation study was carried out. Table IV reports the effect of removing groups of input features using the MAE, and the Concordance Correlation Coefficient (CCC) to quantify such effect. The first column corresponds to the original model, which uses the complete set of features. Values indicating a greater deterioration in the metrics have been boxed. Results show that the most critical sets of features are Collision and Intrusion, which include the collision flags and the social space intrusion variables. Other important features are those in Success, which indicate the task’s completeness. On the contrary, context features in Urgency (task urgency and robot speed) have a limited influence on the output of the model. This can be explained by certain inconsistencies introduced by the LLM, among some context variables. Such inconsistencies are expected to diminish as LLMs continue to improve in accuracy and consistency.
All features Scenario Pose Speed Success Distance Collision Intrusion Near Danger Path Urgency Risk Importance
MAE 0.160 $0.183$ $0.167$ $0.165$ $0.190$ $0.173$ $0.190$ $\boxed{0.191}$ $0.165$ $0.164$ $0.165$ $0.162$ $0.161$ $0.165$
CCC $\mathbf{0.776}$ $0.700$ $0.753$ $0.760$ $0.681$ $0.726$ $\boxed{0.650}$ $0.652$ $0.755$ $0.754$ $0.757$ $0.774$ $0.768$ $0.742$
TABLE IV: Ablation study providing the MAE and CCC values for the proposed model and several ablations. The ablations are as follows: general scenario features -goal thresholds and existence flags- (Scenario); robot pose (Pose); robot speed and acceleration (Speed); success related features -success flag and distance to goal- (Success); distances to humans, objects, and walls (Distance); flags indicating collisions with humans, objects, and walls (Collision); social space intrusions features (Intrusion); features related to the number of near humans (Near); features related to potential danger -minimum TTC, maximum panic, and maximum fear- (Danger); features related to path evolution -path efficiency ratio, step ratio and episode end- (Path); context features related to urgency -task urgency and robot speed- (Urgency); context features related to risk -task risk, distances and justification of bumping into humans and objects- (Risk); and context features related to importance of the task, the human comfort and the robot predictability (Importance).
V-D Qualitative results
To test the model from a qualitative perspective, we generated three simulated scenarios in which people are located frontally between the robot’s initial position and the goal position (Fig. 4(a)). The robot moves at different speeds following several trajectories (a total of $101$) that deviate a variable distance from the straight line to the goal. The top left image in Fig. 4(a) shows a subset of these trajectories as dotted curves. Additionally, four different contexts are considered for each scenario, each with distinct urgency, importance, and risk among other variables: “A robot is working with lab samples. The samples contain a deadly virus” (lab); “A restaurant robot is looking for a fire extinguisher, as it just detected a fire” (fire); “An office assistant robot keeps track of who is in the office today” (office); and “A delivery robot is navigating in a hospital. It works with fragile objects” (fragile).
(a) Robot trajectories and scenarios of Figs. 4(b), 4(c), and 4(d).
@@ V-D Qualitative results
Figure 4: Depiction of the output of the learned metric for different trajectory variations, situations, contexts and speeds. In all scenarios, the robot and the goal are in the same position, as shown in Fig. 4(a), in orange at the bottom and green at the top, respectively. In Figs. 4(b) and 4(d), there is a single human; in the first case static, in the second approaching the robot’s initial location. In Fig. 4(c), there are three humans, as depicted in Fig. 4(a). To reach its goal, the robot follows different trajectories, with varying degrees of divergence from the central line (top-left image of Fig. 4(a)). The context identifiers are shown on the top-left of each figure and described in Sec. V-D. Speed is color-coded:   0.20 m/s   0.40 m/s   0.80 m/s   1.60 m/s.
Figures 4(b), 4(c), and 4(d) show the trained model’s output for these three scenarios. The horizontal axis represents the maximum deviation of the trajectory from the straight line —the sign refers to the side, left or right. The plots show the score assigned by the model to each trajectory. As these figures show, the model learns to distinguish among contexts. The highest speed is penalized more in the contexts labeled as lab and fragile in Fig. 4, where the robot would generally be expected to move slowly due to the risk of the task. In these two contexts, the robot is also expected to maintain a wider distance from humans, which explains why trajectories closer to the human (i.e., with smaller deviations) receive lower scores than in the other contexts. In contrast, the fire context reflects a high-urgency situation: the lowest speed results in scores close to zero on all trajectories and scenarios, and closer proximity to humans is more acceptable.
@@
Finally, notable differences can be observed between the two scenarios with one person (Figs. 4(b) and 4(d)). When the person moves, the robot should move faster than when the person is static. This aligns with the expectation that the robot should actively maneuver in time to avoid the moving person safely. Additionally, deviating farther from the straight line is more penalized than in the static scenario, which can be explained by the fact that, once the robot has safely avoided the human, the distance from them is greater than in the stationary person scenario, as the person moves away from the robot.
Overall, these qualitative results show that the model not only distinguishes among contexts but also captures the balance between speed, proxemics, and urgency, and that it aligns with human expectations of safe and efficient behavior.
@@ VI Conclusions and Future Work
This paper described how All-encompassing Learned Task-wise metrics [francis2023principles] can be learned following a data-driven approach in the context of social robot navigation. The experimental results support the idea that the approach is feasible and that our metric correlates with raters’ scores better than common analytic metrics. The quantitative results achieve MSE and MAE of $0.0457$ and $0.160$, respectively. The qualitative results show that the learned metric is sensitive to both the content of the scenarios and the context of the tasks.
@@
Our envisioned future work spans two different areas: a) increasing the dataset size with new trajectories and trajectories from existing non-rated datasets [thorDataset2019, martin2019jrdb], and b) development of domain-specific deep learning architectures that can more efficiently leverage geometric relationships.